How many ounces of a 15% saline solution must be added to 30 ounces of a 5% saline to make a 10% saline solution?
Found 2 solutions by checkley75, Earlsdon:
Answer by checkley75(3666)

.15x+.05*30=.10(x+30)
.15x+1.5=.10x+3
.15x-.10x=3-1.5
.05x=1.5
x=1.5/.05
x=30 oz. of 15% saline solution is needed.
proof
.15*30+.05*30=.10(30+30)
4.5+1.5=.10(60)
6=6

Answer by Earlsdon(6103)

Let x = the required number of ounces of 15% saline solution.
Change the percentages to their decimal equivalents:
15% = 0.15
5% = 0.05
10% = 0.1
From the problem description, you can write:
(0.15)x + (0.05)(30) = (0.1)(x+30) Simplify and solve for x.
0.15x+1.5 = 0.1x+3 Combining like-terms, you subtract 0.1x from both sides.
0.05x+1.5 = 3 Now subtract 1.5 from both sides.
0.05x = 1.5 Finally, you divide both sides by 0.05
x = 30
You will need to add 30 ounces of a 15% saline solution to 30 ounces of a 5% saline solution to obtain 60 ounces of a 10% saline solution.